OpenStudy (anonymous):

Which equation can be used to solve the following word problem? Seth has 8 more nickels than dimes, and the total value of [NGT]his coins is $2.35. How many nickels does [NGT]he have? Let n represent the number of nickels.

OpenStudy (joshawad):

just to be sure D =8+n right

OpenStudy (anonymous):

.05(n) + .10(n-8) = 2.35 I can not see your choice of answers.

OpenStudy (anonymous):

answers could be 0.10n + 0.05n + 8 = 2.35 0.10n + 0.05(n + 8) = 2.35 0.10(n – 8) + 0.05n = 2.35 0.10(n + 8) + 0.05n = 2.35

OpenStudy (anonymous):

d = n + 8 .10(n + 8) + .05n = 2.35

OpenStudy (anonymous):

0.10(n – 8) + 0.05n = 2.35 is the answere

OpenStudy (anonymous):

are you sure @kelliegirl33

OpenStudy (anonymous):

I messed up...I read it wrong....I thought there were 8 more dimes then nickels. Since there are 8 more nickels then dimes it would be.. d = n - 8 .10d + .05n = 2.35 now sub n - 8 in for d in the 2nd equation .10(n - 8) + .05n = 2.35 sorry about that...coxr15 is correct

OpenStudy (anonymous):

ok thanks to both of you!! @coxr15 @kelliegirl33